Circular No. 6862

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0421+560 AND CI CAMELOPARDALIS R. M. Hjellming and A. J. Mioduszewski, National Radio Astronomy

Observatory, report: “Very Large Array observations of CI Cam on Apr. 3.83 UT reveal the presence of an unresolved point source with preliminary flux densities of 120, 470, 650, 500, and 420 mJy at 1.4, 4.9, 8.4, 14.9, and 22.5 GHz, respectively. This indicates that it was optically thick at the two lower frequencies and close to optically thin at the two higher frequencies. The size of the source is < 0′′.1 at 22.5 GHz. We reanalyzed the 1.4-GHz data for Apr. 2.63 (IAUC 6857) and were able also to image the source at 4.9 and 8.4 GHz, so we can report preliminary flux densities of 20, 150, and 280 mJy at these three frequencies on Apr. 2.63; at that time, the spectrum was optically thick, with a spectral index of ∼ 1.5. The rapid variability of factors of ∼ 6, ∼ 3, and ∼ 2.3 at 1.4, 4.9, and 8.4 GHz, respectively, indicates that this is synchrotron radio emission that is probably related to the x-ray transient XTE J0421+560 (IAUC 6855, 6857). Since the position of the unusual radio source coincides with the ‘symbiotic’ star CI Cam, the radio emission and x-ray transient can both be identified with CI Cam.”

E. L. Robinson, W. F. Welsh, M. T. Adams, and M. E. Cornell, University of Texas and McDonald Observatory, report: “We obtained a wide-field R-band CCD image of the field of XTE J0421+560 on Apr. 2.1 UT with the Prime Focus Camera on the 0.7-m telescope at McDonald Observatory. The peculiar variable star CI Cam (= MWC 84), which is near the center of the XTE error circle and at the position of the radio source found by Hjellming and Mioduszewski, was at R ∼ 7.1. According to Bergner et al. (1995, A.Ap. Suppl. 112, 221), the R magnitude of MWC 84 is normally between 10.47 and 10.69. This optical outburst and the emergence of He II emission lines in its spectrum (Wagner and Starrfield, IAUC 6857) make CI Cam the probably optical counterpart of XTE J0421+560.”

COMET C/1998 F1 (SOHO) C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports the discovery by D. Biesecker of a comet in both C3 and C2 coronagraphic data. The comet was estimated at mag 4, and a tail was clearly visible in the C2 data. Further details with observations and orbit are given on MPEC 1998-G11.

1998 UT  α_{2000}  δ_{2000}
Mar. 21.818  0^{h}15.^{m}9  − 1^{o}08′
1998 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green